a) An organ pipe open at both ends has a radius of 4 cm and a length of 5 m. What is the frequency (in Hz) of the third harmonic?.

b) Calculate the frequency of sound you would hear as you drove at 15m/s toward a sound source emitting a sound of frequency 400 Hz.


1
Expert's answer
2021-09-23T15:48:36-0400

(a) "f=nv/(2L)=3\cdot344/(2\cdot5)=103.2\ (Hz)"


(b) "f=f_0(1+u/c)=400\cdot(1+15/344)=417.4\ (Hz)"